OpenStudy (mony01):

Show that the components of vector in a coordinate system rotated by an angle can be given in terms of the original components as

OpenStudy (mony01):

\[A _{x'}=A _{x}\cos \phi+A _{y}\sin \phi \]

OpenStudy (mony01):

\[A _{y'}=-A _{x}\sin \phi+A _{y}\cos \phi \]